OpenStudy (anonymous):

solve x^2 - 18x + 81= 36

OpenStudy (anonymous):

algebra 2

OpenStudy (anonymous):

@AndrewKaiser333

OpenStudy (andrewkaiser333):

ok x times x = x now it is x-18x+81=36

OpenStudy (andrewkaiser333):

now x-18x = -17x

OpenStudy (anonymous):

I think Im supposed to solve with quadratic formula

OpenStudy (andrewkaiser333):

now we got -17x+81=36 now you - 81 by 36 so 36-81= -45

OpenStudy (andrewkaiser333):

now it is -17x=-45

OpenStudy (andrewkaiser333):

divide -45 by -17 you get - 2.6= x

OpenStudy (andrewkaiser333):

i rounded it

OpenStudy (anonymous):

im supposed to factor or use quadratic formula tho

OpenStudy (andrewkaiser333):

or it would be 2.647058824=x

OpenStudy (andrewkaiser333):

that is how it shows me when i did it and explained it

OpenStudy (anonymous):

can you factor it?

OpenStudy (andrewkaiser333):

i hope it works for ya

OpenStudy (andrewkaiser333):

as a fraction

OpenStudy (andrewkaiser333):

that is hard

OpenStudy (anonymous):

ok thanks I got a square root with quadratic formula

OpenStudy (anonymous):

(x+15(x+3) while factoring

OpenStudy (andrewkaiser333):

sorry i got a headache do not want to go in like that lol

OpenStudy (anonymous):

ok haha thanks